Complete the square:
3y^2+9y = -2
The (2) means squared - don't know how to type it.

Answer by ewatrrr(24785)


Hi
Completing the square:
3y^2+9y = -2 |Note use of ^ (uppercase 6)
3(y^2 + 3y) + 2 = 0
3[(y + 3/2)^2 -9/4] + 2 = 0
3(y + 3/2)^2 - 27/4 + 2 = 0
3(y + 3/2)^2 - 19/4 = 0